OpenStudy (anonymous):

Find F(x+h)-f(x)/h if f(x)= 1/x

hartnn (hartnn):

first find f(x+h) and tell,can u ?

OpenStudy (anonymous):

1/(x+h)?

hartnn (hartnn):

right,so u have\[\frac{\frac{1}{x+h}-\frac{1}{x}}{h}\] right? now cross-multiply.

OpenStudy (anonymous):

How....?

hartnn (hartnn):

\[\frac{\frac{1}{x+h}-\frac{1}{x}}{h}=\frac{x-(x+h)}{xh(x+h)}\] got this? whats numerator simplified to ?

OpenStudy (anonymous):

Its simplified to x, how'd you move the h up.....

OpenStudy (anonymous):

*h

hartnn (hartnn):

i did not move h up,it was previously in the denominator,and still in the denominator. \[\frac{\frac{a}{b}}{c}=\frac{a}{bc}\] and the numerator would be -h. x-x-h=-h

OpenStudy (anonymous):

Ahh okay. So whats the next step?

OpenStudy (anonymous):

and is the a/b/c something I'll learn in calc?

hartnn (hartnn):

do u see something common in num and denom. ? if yes cancel that out and tell me what remains, thats your final answer. that a/b/c is a school stuff,which u should have known.....

OpenStudy (anonymous):

I had a bad teacher for Pre-Calc.

hartnn (hartnn):

so what u got after cancelling?

OpenStudy (anonymous):

x(x+1) ?

OpenStudy (anonymous):

-1/x(x+1)?

hartnn (hartnn):

thats your final answer :)

OpenStudy (anonymous):

Wait I was right?

hartnn (hartnn):

u just missed h -1/x(x+h) is the final answer

OpenStudy (anonymous):

grrr and thanks!

hartnn (hartnn):

welcome :)